Hello Sir,
I want to know "Octal to Decimal conversion"- formula
&
Hexadecimal to Decimal conversion- formula.
Please help me in this.
Thanks

Answer by Edwin McCravy(20055)

I'm assuming you mean converting whole numbers, integers.  
Converting fractions is different.

For octal, the algorithm is this, starting with the first digit

Multiply by 8
Add next digit
Multiply by 8
Add next digit
.....

In hex, just use 16 instead of 8, and remember that A=10, B=11, etc.

until you run out of digits.

Suppose you have the octal number 

647210538

Start with first digit 6

               6
      6*8+4 =  52
     52*8+7 =  423
    423*8+2 =  3386
   3386*8+1 =  27089
  27089*8+0 =  216712
 216712*8+5 =  1733701  
1733701*8+3 = 13869611

647210538 = 1386961110

Suppose we have the hex number:

D59A52C16

You might want to write it this way first:

(13)59(10)52(12)

Start with D which is 13

                 13
      13*16+5  = 213
     213*16+9  = 3417
    3417*16+10 = 54682
   54682*16+5  = 874917
  874917*16+2  = 13998674
13998674*16+12 = 223978796

So D59A52C16 = 22397879610 
 
It's very easy with a TI graphing calculator, because when
you press *16+ after a calculation you don't have to reenter
the last number, it just automatically uses it as Ans and you
see Ans*16+.  So it's very fast on a TI.

Edwin